OpenStudy (anonymous):

Simplify the expression the root of negative one all over the quantity of four minus three i minus the quantity of two plus five i. Really don't get this please help.

OpenStudy (anonymous):

\[\sqrt{-1} \over \left( 4-3i \right) - \left( 2 +5i \right)\]

ganeshie8 (ganeshie8):

simplify the denominator first

ganeshie8 (ganeshie8):

combine like terms..

OpenStudy (anonymous):

\[\sqrt{-1} \over \left( 2-5i \right)\]

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

try again

OpenStudy (anonymous):

is the denominator -8 +15i?

ganeshie8 (ganeshie8):

\(\large \sqrt{-1} \over \left( 4-3i \right) - \left( 2 +5i \right)\) \(\large \sqrt{-1} \over 4-3i - 2 - 5i \) \(\large \sqrt{-1} \over 2-8i \)

ganeshie8 (ganeshie8):

and since \(\sqrt{-1} = i\), ur numerator changes to \(i\)

OpenStudy (anonymous):

oh okay I'm sorry and from there the numerator is i sqrt 1

ganeshie8 (ganeshie8):

\(\large \sqrt{-1} \over \left( 4-3i \right) - \left( 2 +5i \right)\) \(\large \sqrt{-1} \over 4-3i - 2 - 5i \) \(\large \sqrt{-1} \over 2-8i \) \(\large i\over 2-8i \)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yes these are my choices though so I dont knwo where to go from there ):

ganeshie8 (ganeshie8):

you need to get rid of i's in the denominator and put this in standard form

rvc (rvc):

i (2+8i)/4+8) = 2i-8/12 =2(i-4)/12 =i-4/6 this is ur answer i suppose @rosepedal1791

rvc (rvc):

this is in continuation with ganesh's ans

OpenStudy (anonymous):

How can I do that? I'm a bit confused. and that isn't one of my choices, this is a really hard question. I learned about imaginary numbers but not to this extent.

rvc (rvc):

hope u understand

rvc (rvc):

just take the conjugate of denominator @rosepedal1791

OpenStudy (anonymous):

yep, take the the conjugate of denominator

ganeshie8 (ganeshie8):

let me show u in latex

OpenStudy (anonymous):

Okay, thank you. But there is no imaginary number in the denominator only the numerator in your answer. so i +4 / 6 ?

ganeshie8 (ganeshie8):

\(\large i\over 2-8i \) multiply conjugate of denominator, top and bottom : \(\large \frac{i}{ 2-8i }\times \frac{2+8i}{2+8i} \) \(\large \frac{i(2+8i)}{ 2^2+8^2 } \) \(\large \frac{2i - 8}{ 68 } \) \(\large \frac{-4 + i}{ 34 } \)

OpenStudy (anonymous):

Oh okay ! Thank you so much, I had no clue I had to multiply the top and bottom, never learned this before I have no clue why it's on my test. Thank you everyone. Im new so how do I give out medals lol

ganeshie8 (ganeshie8):

:) its kindof like rationalizing denominator...

ganeshie8 (ganeshie8):

nobody likes to leave radicals in the denominator

ganeshie8 (ganeshie8):

so we multiply wid the conjugate and get rid of them

ganeshie8 (ganeshie8):

similarly, we dont want "i"s in the denominator, so we multiply wid conjugate and fix the denominator

OpenStudy (anonymous):

Rationalizing denominators, never got taught that. Teachers can be silly. But thank you I get what you are saying now. i over 2 - 8i didn't look right at first lol (:

ganeshie8 (ganeshie8):

good to hear that :)

rvc (rvc):

keep it up!! @rosepedal1791